OpenStudy (anonymous):

Based on the box-and-whisker plot shown below, match each term with the correct value. Median Range 25th percentile 75th percentile Inter-quartile Range

OpenStudy (anonymous):

median=11

OpenStudy (anonymous):

range=12

OpenStudy (anonymous):

thats all i know sorry

OpenStudy (anonymous):

median = 11 25th percentile (or 1st quartile) = 9 75th percentile (or 3rd quartile) = 14 The range is the difference between the 1st quartile and the 3rd quartile which is 14 - 9 = 5

OpenStudy (anonymous):

any questions ?

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

your welcome...glad to help

OpenStudy (anonymous):

but I still don't know what the inter quartile is??

OpenStudy (anonymous):

the inter quartile is Q3 - Q1........Q3 is also the 75th percentile and Q1 is the 25th percentile

OpenStudy (anonymous):

Ohhhh ok thanks :)

OpenStudy (anonymous):

anything else ?

OpenStudy (anonymous):

are you good with probabilities? :D

OpenStudy (anonymous):

not really.....but I can look at it and see if I can help...what you got ?

OpenStudy (anonymous):

Given the following cards, what is the probability of drawing a blue card, not replacing it, and then drawing another blue card?

OpenStudy (anonymous):

how many cards are we talking about ?

OpenStudy (anonymous):

you have 5 cards.....3 are blue......so the probability of picking a blue card on the first turn is 3/5. Not replacing it would mean there are now 4 cards....and the probability of picking a blue would be 2/4 or 1/2. So the total probability would be 3/5 * 1/2 = 3/10. You might want to get a second opinion on this.....I am not the best at probabilities.

OpenStudy (anonymous):

I got to go now...seriously, you might want to get a second opinion on this :)

OpenStudy (anonymous):

your explanation does make sense and that was one of the answers to chose from so I just went with that. thank you so much.

OpenStudy (anonymous):

your welcome....I just hope it is correct :/